Chemistry Lab
There was little gas and low organic carbon associated with the majority of the cores recovered during Leg 171B. Hydrocarbons were monitored at each site as there were safety concerns for this region. Gas values entered into the JANUS data were retrieved easily to monitor trends. Few problems were associated with the equipment, though new Gas Chromatograph (GC) Chemstation software was found to be incompatible with the valve switcher. A new version of the software is expected to work as anticipated. There was also a problem with a Coulometer application in LabView that will require more work. Replacement parts and valves for the Atomic Absorption's (AA) exterior acetylene manifold were ordered for port call service. Other maintenance in the lab was routine.

Computer Services
Many related tasks were performed with the implementation of the JANUS database on this leg. There were hardware upgrades, an OS upgrade, the initialization of the database, user training for the scientists and specialists, bugs to fix, user resistance to overcome, etc. Full user support was available from Tracor's Paul Albright and Glen Corser for the JANUS applications, and they continued working on the Business Objects Users software queries throughout the leg. The deployment of JANUS on Leg 171 was a success; it is basically up and functioning as intended. There were problems mirroring data across the hard drives and implementing DecSafe, an application that switches computers if a processor fails. Those shortcomings will soon be addressed.

The latest version of AppleCORE was used successfully by the scientists to prepare barrel sheets. It was found to be flexible and allowed an almost unlimited amount of information to be entered. A VIDEO program was sent out to replace the old corelog video display. It was modified to some degree to change sort orders, precision, and to add scrolling. On several occasions at the end of the leg the maximum number of users on the Grosbeck/USERVOL server was reached. A license change is planned. There were several problems with cc:mail, aggravated by the holidays at Texas A&M;University (TAMU). Bulletin boards failed to propagate for one period of time. There was a crash of the system shortly after leaving Barbados when new addresses were being added. The last short run of the FDDI fiber optics cable was run. These will be terminated with the rest of the system cables during the Charleston port call.

Core Lab
During Leg 171A's LWD effort, SUN workstations were set up in the aft end of the core lab to support the scientists reviewing the data collected. Working in the area during daylight hours was difficult as the entire deck above was being chipped, scaled, and then repainted. Some people used hearing protection, whereas others elected to work later hours.

There were few problems accommodating the recovered sediments studied on Leg 171B. Drilling breaks and logging activities provided time for the special requirements that support K/T boundary cutting and sampling efforts.

As this was the first leg for a fully implemented JANUS database, double entry into S1032 corelog was made as a backup. During the last sites, maintaining the core entry white board was dropped, as most of the drilling information is now available on the workstations that were logged onto JANUS. The white board is still handy for planning and scheduling Adara tool deployments, special sampling, and paleontological ages.

Core recovery was quite high from the relatively shallow-water sites during this five-week effort, but chert and harder layers encountered precluded achieving some of the higher estimates. Rotary coring and the logging program provided the time necessary for the lab to catch up and conduct sample parties.

Several critical boundaries, including the K/T boundary, were recovered, and many cutting and sampling suggestions and ideas were evaluated. Some special tools, now stored in a K/T kit, were used. Thin aluminum oxide cutting disks were tried on the rock saw with marginal success. They were too fragile for much regular use and they ablated material off onto the core that could not easily be cleaned off. Cores with organic-rich layers were recovered and handled in a special manner. They were placed in foil bags and purged with nitrogen to preserve them. They will be held until a sample plan is worked out. Some cores contained colorful laminated zones and were selected to join the K/T boundary cores as a part of an exhibit supporting port call Public Relation efforts. All of these sections will be shipped to the Bremen Core Repository for future work and storage. A second curatorial person was a fine complement to the staff, allowing full-time help and oversight in the lab and catwalk.

Magnetics Lab
This was the first heavy recovery leg for the new cryogenic magnetometer; both the hardware and the software performed satisfactorily. There were a few problems to puzzle over, but the extra sensitivity was appreciated. Core sections contaminated with metal bits possibly contributed to some unexplained flux jumps and slow recovery of stability after the chill water was off for 2 hr. While the data was modified by JANUS to add depths, the raw data files were stored on DATA1 file server.

The Tensor orientation tools were used regularly. The data retrieval program was modified to run on a Pentium laptop computer. One of the tools is being returned with a damaged battery contact and intermittent problems.

Microscope/Photography Labs
The microscopes were set up to meet individual preferences, and time was spent finding special equipment and supplies to fill requests. Occasionally, we were asked to change lenses or readjust the optics or illumination and replace bulbs. Notes were taken for a port call microscope service call so problems can be directly addressed and also for some instruction on aligning one new model microscope.

Close-up photographs of some of the critical boundaries were taken before and after the surfaces were scraped clean. Duplicate slides of the K/T boundary were taken so each of the scientists could have one, thereby reducing the number of times the boundary would be handled. Macrophotographs of some ammonite fossils were taken. No microphotographs were requested. There were no problems with the lab equipment.

Paleontology Lab
The facilities were fully utilized with seven paleontologists sharing the space. Once all requests were satisfied during the initial setup, no problems were reported. A situation attributed to gray water fumes was reported in the aft end of the microscope lab. Better termination of an old drain line seemed to fix the problem. The chemists responded to most of the consumable requests during the leg.

Physical Properties
The physical properties scientists were familiar with the instruments and had few problems. Those making thermal conductivity measurements found less scatter in the values with the newer TKO4, so it was preferred over the WHOI multiprobe device. The users found that core flow was not affected using the TKO4's single needle probe. Sets of Adara downhole temperature measurements were collected at two drill sites.

The storekeeper travelled early to Panama to meet the ship's agent, to familiarize himself with the area, and to help locate the arriving equipment. Hotel reservations were verified, and outgoing travel arrangements verified. The port call in Panama went well once shipments started, and everything was received in good order.

During Leg 171A, time was spent with the engineer who joined the staff to study migrating the MATMAN S1032 database into an interim Foxpro database (the software the engineering group is accustomed to using). It had been discussed that the storekeeper would become involved with the engineering inventories, but there was no involvement this leg.

There was time to make a physical count of several of the storage areas, with Hold Stores (HS) getting the most attention. On-hand numbers were adjusted and orders made accordingly. Two pieces of laboratory equipment, the HP FAXATRON X-radiograph and the Spectrex PC-2000 particle analyzer, were included in the shipment to ODP. Core boxes will be sent to the Bremen Core Repository with trans-shipped supplies from the Gulf Coast Repository.

The opportunity of sailing on the LWD part of Leg 171 afforded time for a JANUS mandated conversion of operating systems on the underway SUN workstations to Solaris 2.5. A UNIX specialist sailed to help with the hardware and software updates and to rewrite the applications as necessary for them to run under the new operating system. There was also the opportunity to write a new SiteFix program for the SUNs and for the underway geophysical specialist to be introduced to nuances of the seismic processing applications. A new a2d trigger was designed and built to replace the external SUN trigger, which did not work with the new version of a2d software.

Both the port and starboard seismic bundles were refurbished and signal leads replaced. The Teledyne single-channel streamer was used during the last leg and was reported to be noisy. A bad leak was located and repaired temporarily, the streamer connectors were cleaned, and the streamers were filled with oil. One magnetometer sensor was leaking fluid and the other sensor was noisy, so both sensors were serviced. A selector switch was added to a front panel in the lab that will allow easier comparisons between the two sensors.

X-ray Lab
The lab was used sparingly during Leg 171B with one scientist making the most use of it and the XRD to identify clay minerals in 45 samples. A few sediment samples were analyzed with the XRF. Some old or seldom-run XRF standards were re-analyzed for comparison with the original numbers. A routine preventative maintenance service call was scheduled for Charleston. Most time and effort was given supporting core recovery and the core lab routine.

To 171B Electronics Support, Special Projects, Safety and Problems

To 171B Table of Contents

Publications Homepage

ODP Homepage